OpenStudy (anonymous):

i keep getting 20x_>-1 but its not the right answer please help the question is Solve five minus square root of the quantity twenty x plus four is greater than or equal to negative three.

OpenStudy (mathmate):

Like this? \[5-\sqrt{20x+4}\ge -3\] Start with isolating x.by moving the radical to the right, numbers to the left, and swapping left and right: \[\sqrt{20x+4}\le 8\] Square both sides, and solve for x. Important to remember that since the domain of sqrt(20x+4) is [-1/5,inf), the condition must be specified in the solution.